Actions
Ops/Comm Scheduling meeting¶
Location : CR11 in R3 @ 15:30 pm on Tuesday 12th
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda¶
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information (Janusz)
- DC report
- BPS Test
- Decay Solenoid status : A_Comment_on_the_Recent_Stability_of_the_Decay_Solenoid.pdf
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/5892/Mice_Hall_Plan-16-03-16.xlsx
- Water system work
- Bellows re-installation and realignment work
- QP/QD work
- Anything else
- April 14th tour : Just a reminder about a series of tours from 14:00 on the afternoon of April 14th.
- Tour schedule : HighPowerTargets_SuggestedProgramme_14Apr2016.xlsx
- Guides : Currently : Pierrick, Victoria, Ken
- April 19th : Outreach visit by school students to paint the wall.
- AOB
- DONM : 19/4/2016
Minutes¶
Present: Mark, Mike, Vicky, Henry, Chris, Steve, Colin, Josef, John, Ed, Ken
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information : stands
- DC report - activities in the Hall this week
- BPS test - Went well after test parameters were prepared. Live drop for Bryan Jones was not possible as ISIS was off but "dead" drop was carried out instead and passed ok. Mark Arnold witnessed frame interlock tests.
- DS - DS presently at 0Amps. Checked feedback of actuators in DSA area and all were fine. Mark and Mike still fine tuning the system.
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/5892/Mice_Hall_Plan-16-03-16.xlsx
- Water system - We are getting chiller that has capacity for 2xSS but there is some doubt if this has 23KW spare for the FC as well. Chiller to be located on gas bottle rack outside hall. Beware FC cooling could be interrupted by forthcoming works on the roof chiller.
- Bellows installtion - Bellows received on Monday and cleaned up. Require approx ~2wks of drilling through for clamping bolt holes.
- QP/QD work - Further installation work by DL this week. Commissioning on schedule for early May
- April 14th tour : MICE tour timings and guides in place. All procedural issues dealt with by ISIS contact
- April 19th : Outreach visit by school students to paint the wall. Students are 16yrs+, Ken states that all risks are taken care of and that hall will need minimal preparation.
- AOB
- Water detector is showing there is water in the trench drain. Has been showing this status for ~2yrs. Needs investigating - Action Pierrick & Craig
- DS readback current is inaccurate requires calibration - Action DL. Ensure cooling channel magnet readback is accurate. Follow Rogers MICE note 474 procedure. - Action Jaroslaw
- Issues arising from last data taking run need to be identified and actioned - Action Boyd
- Internal MICE safety meeting requested to discuss PRY meshing and forthcoming safety procedures relating to RF and LH2 systems - Nichols
Updated by MacWaters, Craig over 7 years ago ยท 11 revisions